# A First-Order Gradient Approach for the Connectivity Analysis of Markov Chains

### v2

@@ IV First-order derivatives and approximations
where $\Delta$ is a perturbation vector such that $\Delta\sim\{-1,1\}^{d}$ uniformly distributed. We can find a stationary point of $J$ using the following basic stochastic approximation recursion initiated at some initial point $x(0)$,
@@
As we show in Appendix -A, standard SPSA is not applicable for solving problem (16), since the constraints imposing $\mathcal{X}$ are hard constraints and therefore cannot be violated. Therefore, we adapt the generic SPSA algorithm in (23) to the constraint set $\mathcal{X}$. First, we derive in Section IV-A an analytical expression for the steepest feasible descent on $\mathcal{X}^{\mathrm{eq}}$. Then, we introduce in Section IV-B a SPSA using this steepest feasible descent, which implies that $x\pm\eta\Delta\in\mathcal{X}^{\mathrm{eq}}$. Subsequently, we present in Section IV-C a modified SPSA algorithm that guarantees that the perturbed values $x\pm\eta\Delta$ stay inside $\mathcal{X}^{\mathrm{pos}}$. Integrating both modifications, the resulting overall SPSA algorithm guarantees that $x\pm\eta\Delta$ is feasible for a feasible $x$, see Section IV-D. Finally, we discuss how to apply SPSA for the extended problem (14) in Section IV-E.
@@ IV-A Analytical derivatives
IV-A1 Directional derivatives on $\mathcal{X}$
Denote the set of all feasible directions leaving $x\in\mathcal{X}$ as
@@ IV-A1 Directional derivatives on $\mathcal{X}$
Then, we can define the directional derivative for continuously differentiable function $J(x)$ in the feasible direction $\delta\in\mathcal{D}(x;\mathcal{X})$ as
@@
Let $x\in\mathcal{X}$ such that $P(x)\in\mathcal{P}$ and take $\delta\in\mathcal{D}(x)$. The directional derivative of $\Pi$ with respect to $\delta$ is given by
@@
IV-A2 Steepest descent direction on $\mathcal{X}_{\varepsilon}$
@@
$\delta(x)=\operatorname*{arg\,min}_{\delta\in\mathcal{D}\left(x;\mathcal{X}\right)}\ \partial_{\delta}J(x),$ (24)
for inner points of $\mathcal{X}$. For any $x\in\mathcal{X}_{\varepsilon}$, $x$ is strictly feasible with respect to non-negativity constraints of $\mathcal{X}$, i.e., $x>0$. Therefore, a feasible direction $\delta\in\mathcal{D}(x;\mathcal{X})$ preserves that $P(x+\eta\delta)\in\mathcal{P}$, for $\eta$ sufficiently small. We can derive the necessary conditions for $\delta$ by writing the equality constraints in $\mathcal{X}$ as a system of equations $Ax=b$, where
$A^{N\times|E|}=\begin{bmatrix}\bar{1}^{\top}_{|V_{1}|}&&\\ &\ddots&\\ &&\bar{1}^{\top}_{|V_{N}|}\end{bmatrix},\text{ and }b=\bar{1}.$ (25)
@@
where $U\in\mathbb{R}^{N\times N}$ is a real orthogonal matrix, $\Sigma$ a diagonal matrix with singular values of $A$ on its diagonal in descending order and $V^{\top}\in\mathbb{R}^{|E|\times|E|}$ also is a real orthogonal matrix. Then, $B$ corresponds to the last $|E|-r$ columns of $V$.
@@
for all $x\in\mathcal{X}_{\varepsilon}$ and normalizing the direction to unit length yields $\delta(x)=\delta^{\prime}(x)/\sum_{\ell\in E}\delta_{\ell}^{\prime}(x)$.
@@ IV-C Keeping SPSA on $\mathcal{X}^{\mathrm{pos}}$
In the previous subsection, we have shown how the equality constraint can be incorporated via a base transformation and a steepest feasible descent direction staying on $\mathcal{X}^{\mathrm{eq}}$ can be found. In the following, we illustrate how to adjust SPSA so that $x+\eta\Delta\in\mathcal{X}^{\mathrm{pos}}$. For this, we adapt Algorithm (23) as follows
@@
As shown by [26], one could apply a constrained SPSA on the original set ${\mathcal{X}}^{\mathrm{pos}}$. To that end, introduce $\{\varepsilon(k)\}\to 0$ for $k\to\infty$, so that ${\mathcal{X}}^{(\mathrm{pos},\varepsilon(k))}\to{\mathcal{X}}^{\mathrm{pos}},\text{ for }k\to\infty.$ Then, Algorithm (23) can be adapted as follows
@@ IV-D The overall SPSA algorithm
In the following, we will solve (16) over $\mathcal{X}_{\varepsilon}$ instead of $\mathcal{X}$, using the first-order algorithm akin (23):
@@ IV-E Constraining the stationary distribution
$\begin{bmatrix}A^{\hat{\pi}}&b^{\hat{\pi}}\end{bmatrix}=\text{rref}\left(\begin{bmatrix}A_{1}&\bar{1}\\ A_{2}&\hat{\pi}\end{bmatrix}\right),$ (35)
@@
$x(k+1)=\text{Proj}_{\mathcal{X}_{\varepsilon}^{\hat{\hat{\pi}}}}\left[x(k)+\alpha(k)\tilde{\delta}_{k}(x(k))\right],$ (36)
Unlike in the previous setting, there is no straightforward closed-form projection on $\text{Proj}_{\mathcal{X_{\varepsilon}^{\hat{\pi}}}}[\cdot]$. However, we can use Dykstra’s projection method of iteratively projecting on convex sets, see [28], provided these projections can be carried out quickly. To find $\text{Proj}_{\mathcal{X}_{\varepsilon}^{\hat{\pi}}}[x]$, let $\mathcal{X}_{\varepsilon}^{\hat{\pi}}=\mathcal{X}^{1}\cap\mathcal{X}^{2}\neq\emptyset$, with
@@ V-A Redistribution function
We now introduce a framework to analyze the connectivity of a Markov chain on a graph whose vertex set is fixed, but the edge set is random. We assume that only a subset of risky edges $\tilde{E}\subset E$ are affected by uncertainty. Any risky edge $(i,j)\in\tilde{E}$ fails (and therefore becomes inaccessible) with probability $q_{ij}\in[0,1]$. Furthermore, we assume that $G(V,E\setminus\tilde{E})$ is strongly connected. If a risky edge $(i,j)$ fails, it means that the original probability mass $P_{ij}$ must be redistributed over the remaining elements in the $i$th row of $P$. Given a realization of accessible edges $\mathcal{E}\subseteq\tilde{E}$, we calculate the adjusted Markov chain $Q(P,\mathcal{E})=(Q_{ij}(P,\mathcal{E}))_{(i,j)\in V^{2}}$, where
@@
$\min_{x\in\mathcal{X}^{(\mathrm{sym},1)}}$ $\mathbb{E}_{\mathcal{E}}\left[S(Q(x,\mathcal{E}),C)\right]$ (43)
where $Q(x,\mathcal{E})=Q(P(x),\mathcal{E})$. It follows that the reversibility condition (R) is satisfied for $P(x)$ via (3) for all strongly connected $(G,x)$ Moreover, let reciprocal edges fail concurrently. Then, (43) is a convex problem. Additionally, if $C$ is a matrix with identical entries $C_{ij}=c>0$, (43) is strictly convex.
@@
$\min_{x\in\mathcal{X}^{(\mathrm{sym},1)}}$ $\sum_{l=1}^{2^{|\mathcal{E}|}}\text{Prob}(\mathcal{E}_{l})\cdot S(Q(x,\mathcal{E}_{l}),C)$ (44)
@@ V-B SPSA for Markov chains on random graphs
We now show how SPSA is easily adapted to handle the random support setting, that is, we solve the stochastic optimization problem
$\min_{x\in{\cal\mathcal{X}}_{\varepsilon}}\quad$ $\mathbb{E}_{\mathcal{E}}\left[S(Q(x,\mathcal{E}),C)\right],$ (47)
where it is assumed that the distribution of $\mathcal{E}$ is unknown.
@@
using graph realizations $\bm{\mathcal{E}}$. To that end, first, observe that
@@
which follows from the independence between observations $\bm{\mathcal{E}}$, chosen $\delta$, and the finiteness of the collection of possible graph realizations. It then follows that
@@
Given a finite number of graph realizations $\bm{\mathcal{E}}$, define
@@ VI Numerical experiments
Consider the grid with bidirectional edges in Fig. 4. The edges in $\tilde{E}$ that can fail (with probability $0.1$) are displayed in red and labeled. It can be easily verified that $G=(V,E\setminus\tilde{E})$ is strongly connected. Our goal is to find a reversible policy $P(x)$ minimizing $\mathbb{E}_{\mathcal{E}}\left[S(Q(x,\mathcal{E}),C^{\hat{\pi}})\right]$, such that its stationary distribution is uniform, i.e., $\pi(P(x))=\frac{1}{N}$. Note that condition (R) implies that $P(x)=P(x)^{\top}$ and so we obtain the following problem:
$\min_{x\in{\cal X}^{\mathrm{sym}}\cap{\cal X}_{\varepsilon}}\quad$ $\mathbb{E}_{\mathcal{E}}\left[S(Q(x,\mathcal{E}),C^{\hat{\pi}})\right],$ (54)
@@
To compare our SPSA approach, we formulate an equivalent problem formulation of minimizing $\mathbb{E}_{\mathcal{E}}\left[S(Q(x(u),\mathcal{E}),C^{\hat{\pi}})\right]$, akin to (33). Then, we solve the problem using the IPOPT solver with a tolerance of $10^{-1}$. To start the IPOPT, we first collect a sample of 10 initial starting points and choose the one with the lowest objective value. We provide the same initial starting point to both SPSA and IPOPT methods and terminate the SPSA algorithm if its objective value improves the IPOPT optimal solution, which we denote by $x^{\text{IPOPT}}$. We evaluate the objective every $l=50.000$ iterations by means of Polyak-Rupert averaging [38, 39], that is, we consider an average of the previous 50% of observations;
@@
and we terminate the algorithm if that solution improves the IPOPT solution, i.e., $|J(\bar{x}_{i}^{(l)})-J(x^{\text{IPOPT}})|<10^{-3}$, for $J(x)=\mathbb{E}_{\mathcal{E}}\left[S(Q(x,\mathcal{E}),C^{\hat{\pi}})\right]$, or when a large number $K=5\cdot 10^{6}$ of iterations is achieved. However, in all our experiments, convergence was reached in less than $K$ steps.
We calibrate SPSA according to (30) with the settings as in Table I and repeat the simulation for $|\tilde{E}|=0,1,\ldots,15$, where $\tilde{E}$ is the set of edges with the lowest labels included in Fig. 4. We show the running time for IPOPT and the SPSA mean and range for 10 repetitions for each $|\tilde{E}|=0,1,\ldots,15$ in Fig. 5. Note that the situation $|\tilde{E}|=0$ corresponds to the fixed support case in which IPOPT is faster. However, our analysis reveals that the running time of IPOPT demonstrates a pattern closely resembling exponential growth. In contrast, SPSA is not subject to a similar exponential increase in computational demand. Specifically, SPSA tends to underperform in smaller instances, which may be attributed to inadequate tuning of hyperparameters. However, it should be noted that the running time of SPSA consistently remains below 1.2 hours across all tested instances. A comparison of SPSA with other solvers, such as the Sequential Least Squares Quadratic Programming (SLSQP) solver from the Scipy package, gave similar results as those presented in Fig. 5.
@@ VII Application: Robotic Surveillance
In this section, we validate the applicability of our method in the context of surveillance, in which problem (14) has a natural interpretation. Here, $\hat{\pi}_{i}$ is the anticipated probability that an intruder appears at a node $i$ (e.g., uniform or based on historical data), so that $P(x)$ with $\pi(P(x))=\hat{\pi}$ that minimizes $S(P(x),C^{\hat{\pi}})$ (or $\mathbb{E}_{\mathcal{E}}\left[S(Q(x,\mathcal{E}),C^{\hat{\pi}})\right]$ in the random support setting), is a surveillance policy with a minimal average number of steps to capture the intruder. Specifically, this surveillance problem with $C=C^{\hat{\pi}}$ has been studied by [11], in which the authors find the optimal surveillance policy with minimal Kemeny constant assuming reversibility of $P(x)$ in the fixed support setting.
@@ VII-1 Fixed support
$\begin{array}[]{rll}\displaystyle\min_{x\in\mathcal{X}_{\varepsilon}^{\hat{\pi}}}&S\left(P(x),C^{\hat{\pi}}\right),\end{array}$ (55)
@@
We compare our non-reversible solution $P(x^{\text{non-rev}})$ with the Markov chain with minimal Kemeny constant assuming reversibility $P(x^{\text{rev}})$ [11]. We report the results for 500 simulations in Table II, for each of the three solutions. As in the experiment by [11], we see that the number of intruders caught increases with a decrease in the Kemeny constant of the surveillance policy; the policy with the lowest Kemeny constant catches the most intruders. Inspecting the $P(x^{\text{non-rev}})$ policy in Fig. 6, we see that a strong level of (clockwise) directionality is achieved (note that the Hamiltonian cycle is the extreme case of directionality), which the $P(x^{\text{rev}})$ policy cannot obtain. This directionality feature is crucial for our approach to obtain lower MFPTs and, thus, higher intruder detection rates.
@@ VII-2 Random support
$\begin{array}[]{rll}\displaystyle\min_{x\in\mathcal{X}_{\varepsilon}^{\hat{\pi}}}&\mathbb{E}_{\mathcal{E}}\left[S\left(Q(x,\mathcal{E}),C^{\hat{\pi}}\right)\right],\end{array}$ (56)
@@ VIII Conclusion and future work
To illustrate why standard SPSA does not work, let us consider the following Markov chain on the complete graph $G$ with $N=3$ nodes and stochastic matrix
@@
$\begin{split}&\text{Proj}_{\mathcal{X}^{1}}[x]=\\ &\,=B^{\hat{\pi}}\left((B^{\hat{\pi}})^{\top}B^{\hat{\pi}}\right)^{-1}(B^{\hat{\pi}})^{\top}\left(x-(A^{\hat{\pi}})^{\dagger}b^{\hat{\pi}}\right)+(A^{\hat{\pi}})^{\dagger}b^{\hat{\pi}},\\ &\,=B^{\hat{\pi}}(B^{\hat{\pi}})^{\top}\left(x-(A^{\hat{\pi}})^{\dagger}b^{\hat{\pi}}\right)+(A^{\hat{\pi}})^{\dagger}b^{\hat{\pi}},\end{split}$
@@
Observe that if the weighted graph $(G,x)$, where $x\in\mathcal{X}^{\mathrm{sym}}$, is strongly connected, the Markov chain $P(x)$ via (3) satisfies (R), i.e., $P(x)\in\mathcal{P}^{\mathrm{rev}}$ [10]. Moreover, it has stationary distribution
@@
This uniqueness implies that $x$ is a unique representation for $P$ via $P(x)$. Moreover, since $x$ is strongly connected as follows from the irreducibility of $P\in\mathcal{P}^{\mathrm{rev}}$, the mapping $P(x)$ is a bijection from $\mathcal{X}^{(\mathrm{sym},1)}\cap\{x\in\mathcal{X}^{\mathrm{sym}}:(G,x)\text{is strongly connected}\}$ to $\mathcal{P}^{\mathrm{rev}}$.
@@
The optimization problem (18) has been studied in detailed in [43], in which the authors proved that the optimal solution $\hat{P}$ is a Hamiltonian cycle. The same paper also shows that $\sum_{i\neq j}M_{ij}=\frac{N(N-1)}{2}$ for all $i$, from which it easily follows that
@@
Consider the family of complete undirected graphs $(G,x)$ with $x\in\mathcal{X}^{\mathrm{sym}}$. The complete graph with $N$ nodes, with $x_{\ell}=1$, for all $\ell\in E$, has been shown to have total effective graph resistance $R_{tot}(x)=N-1$, see [6]. Using this fact in combination with (9) and (10) provides the lower bound in (21), that is
@@
The upper bound in (21) follows from the fact that the Hamiltonian graph with maximal $R_{tot}(x)$ is a circular graph (as adding edges to $G$ cannot increase $R_{tot}(x)$, [10]). In [44], it has been shown that for a circular graph $(G,x)$ of size $N$, $R_{tot}(x)=\frac{N^{3}-N}{12}$, where $x_{\ell}=1$ if $\ell\in E$ is part of the circle and $x_{\ell}=0$ otherwise. Since $(G,x)$ is a circular graph, we have $\sum_{\ell\in E}x_{\ell}=N$. Rewriting (9) yields:

### v4

@@ IV First-order derivatives and approximations
where $\Delta\sim\text{Rad}^{d}$ is a perturbation vector with entries independently and uniformly distributed over $\{-1,1\}$. We can find a stationary point of $J$ using the following basic stochastic approximation recursion initiated at some initial point $x(0)$,
@@
As we show in Appendix -A, standard SPSA is not applicable for solving problem (15), since the constraints imposing $\mathcal{X}$ are hard constraints and therefore cannot be violated. Therefore, we adapt the generic SPSA algorithm in (22) to operate within the constraint set $\mathcal{X}$. To that end, we derive in Section IV-A an analytical expression for the steepest feasible descent on $\mathcal{X}^{\mathrm{eq}}$. Then, in Section IV-B, we introduce a version of SPSA that estimates this descent direction, ensuring that $x\pm\eta\Delta\in\mathcal{X}^{\mathrm{eq}}$. Independently, Section IV-C presents an SPSA variant following [14] that is designed to maintain the feasibility of the perturbed values $x\pm\eta\Delta$ with respect to the inequality-constrained set $\mathcal{X}^{\mathrm{pos}}$ across all iterates of $x$. To ensure this, we require that each iterate $x$ is projected back onto the restricted set $\mathcal{X}^{(\mathrm{pos},\varepsilon)}$, so that the resulting perturbations remain within $\mathcal{X}^{\mathrm{pos}}$. In Section IV-D, we then combine these two components, with equality feasibility from Section IV-B and inequality feasibility from Section IV-C, into a SPSA algorithm that guarantees $x\pm\eta\Delta$ is feasible for any $x\in\mathcal{X}_{\varepsilon}$, thereby solving problem (15) over $\mathcal{X}_{\varepsilon}$. Lastly, we discuss how to apply SPSA to the extended problem (16) in Section IV-E, where a similar approach is used to optimize over $\mathcal{X}^{\hat{\pi}}_{\varepsilon}$.
@@ IV-A Analytical derivatives
IV-A1 Directional derivatives
Denote the set of all feasible directions leaving $x\in\mathcal{X}_{\varepsilon}$ with respect to the set $\mathcal{X}$ as
@@ IV-A1 Directional derivatives
Then, we can define the directional derivative for a continuously differentiable function $J(x)$ in the feasible direction $\delta\in\mathcal{D}(x;\mathcal{X})$ as
@@
Let $x\in\mathcal{X}_{\varepsilon}$ such that $P(x)\in\mathcal{P}$ and take $\delta\in\mathcal{D}(x,\mathcal{X})$. The directional derivative of $\Pi$ with respect to $\delta$ is given by
@@
IV-A2 Steepest descent direction
@@
$\delta(x)=\operatorname*{arg\,min}_{\delta\in\mathcal{D}\left(x;\mathcal{X}\right)}\ \partial_{\delta}J(x).$ (23)
For any $\in\mathcal{X}_{\varepsilon}, xx$ is strictly feasible with respect to non-negativity constraints of $\mathcal{X}$, i.e., $x>0$. Therefore, a feasible direction $\delta\in\mathcal{D}(x;\mathcal{X})$ preserves that $P(x+\eta\delta)\in\mathcal{P}$, for $\eta$ sufficiently small. We can derive the necessary conditions for $\delta$ by writing the equality constraints in $\mathcal{X}$ as a system of equations $Ax=b$, where
$A^{N\times|E|}=\begin{bmatrix}\bar{1}^{\top}_{|E_{1}|}&&\\ &\ddots&\\ &&\bar{1}^{\top}_{|E_{N}|}\end{bmatrix},\text{ and }b=\bar{1}.$ (24)
@@
where $U\in\mathbb{R}^{N\times N}$ is a real orthogonal matrix, $\Sigma$ a diagonal matrix with singular values of $A$ on its diagonal in descending order and $V^{\top}\in\mathbb{R}^{|E|\times|E|}$ also is a real orthogonal matrix. Then, $B$ corresponds to the last |E|r|E|-r columns of $V$.
@@
for all $x\in\mathcal{X}_{\varepsilon}$ and normalizing the direction to unit length yields $\delta(x)=\delta^{\prime}(x)/\|\delta^{\prime}(x)\|_{2}$.
@@ IV-C Keeping SPSA on $\mathcal{X}^{\mathrm{pos}}$
In the previous subsection, we have shown how the equality constraint can be incorporated via a base transformation and a steepest feasible descent direction staying on $\mathcal{X}^{\mathrm{eq}}$ can be found. In the following, we illustrate how to adjust the SPSA algorithm in (22), so that $x+\eta\Delta\in\mathcal{X}^{\mathrm{pos}}$:
@@
As shown by [14], one could apply a constrained SPSA on the original set ${\mathcal{X}}^{\mathrm{pos}}$. To that end, introduce $\{\varepsilon(k)\}\to 0$ for $k\to\infty$, so that ${\mathcal{X}}^{(\mathrm{pos},\varepsilon(k))}\to{\mathcal{X}}^{\mathrm{pos}},\text{ for }k\to\infty.$ Then, the algorithm in (22) can be adapted as follows
@@ IV-D The overall SPSA algorithm
By integrating the approaches from sections IV-B and IV-C, we obtain the following SPSA algorithm:
@@ IV-E Constraining the stationary distribution
[…]
@@
$x(k+1)=\text{Proj}_{\mathcal{X}_{\varepsilon}^{\hat{\hat{\pi}}}}\left[x(k)+\alpha(k)\tilde{\delta}_{k}(x(k))\right].$ (35)
Unlike in the previous setting, there is no straightforward closed-form projection on $\text{Proj}_{\mathcal{X_{\varepsilon}^{\hat{\pi}}}}$$\cdot$]. However, we can use Dykstra’s projection method of iteratively projecting on convex sets, see [29], provided these projections can be carried out quickly. To find $\text{Proj}_{\mathcal{X}_{\varepsilon}^{\hat{\pi}}}[x]$, let $\mathcal{X}_{\varepsilon}^{\hat{\pi}}=\mathcal{X}^{1}\cap\mathcal{X}^{2}\neq\emptyset$, with
@@ V-A Redistribution function
[…]
@@
$\min_{x\in\mathcal{X}^{(\mathrm{sym},1)}}$ $\mathbb{E}_{\mathcal{E}}\left[S(Q(x,\mathcal{E}),C)\right],$ (42)
where $Q(x,\mathcal{E})=Q(P(x),\mathcal{E})$. It follows that the reversibility condition (R) is satisfied for $P(x)$ via (3) for all strongly connected $(\mathscr{G},x)$ Moreover, let reciprocal edges fail concurrently. Then, (42) is a convex problem. Additionally, if $C$ is a matrix with identical entries $C_{ij}=c>0$, (42) is strictly convex.
@@
$\min_{x\in\mathcal{X}^{(\mathrm{sym},1)}}$ $\sum_{l=1}^{2^{|\mathcal{E}|}}\text{Prob}(\mathcal{E}_{l})\cdot S(Q(x,\mathcal{E}_{l}),C).$ (43)
@@
Similarly to Section IV, we solve these problems over $\mathcal{X}_{\varepsilon}$ and $\mathcal{X}_{\varepsilon}^{\hat{\pi}}$, respectively, which ensures that the SPSA algorithm remains applicable under the given constraints.
@@ V-B SPSA for Markov chains on random graphs
We now show how SPSA is easily adapted to handle the random support setting, that is, we solve the stochastic optimization problem (44), where it is assumed that the distribution of $\mathcal{E}$ is unknown.
@@
[…]
@@
[…]
@@
[…]
@@ VI Numerical experiments
[…]
@@
[…]
@@
[…]
We calibrate SPSA according to (29) with the settings as in Table I and repeat the simulation for $|\tilde{E}|=0,1,\ldots,15$, where $\tilde{E}$ is the set of edges with the lowest labels included in Fig. 4. We show the running time for IPOPT and the SPSA mean and range for 10 repetitions for each $|\tilde{E}|=0,1,\ldots,15$ in Fig. 5. Note that the situation $|\tilde{E}|=0$ corresponds to the fixed support case in which IPOPT is faster. However, our analysis reveals that the running time of IPOPT demonstrates a pattern closely resembling exponential growth. In contrast, SPSA is not subject to a similar exponential increase in computational demand. Specifically, SPSA tends to underperform in smaller instances, which may be attributed to inadequate tuning of hyperparameters. Nevertheless, it should be noted that the running time of SPSA consistently remains below 1.2 hours across all tested instances. A comparison of SPSA with other solvers, such as the Sequential Least Squares Quadratic Programming (SLSQP) solver from the Scipy package, gave similar results to those presented in Fig. 5.
@@ VII Application: robotic surveillance
[…]
@@ VII-1 Fixed support
[…]
@@
We compare our non-reversible solution $P(x^{\text{non-rev}})$ with the Markov chain with minimal Kemeny constant assuming reversibility $P(x^{\text{rev}})$ [11]. We report the results for 500 simulations in Table II, for each of the two solutions. As in the experiment by [11], we see that the number of intruders caught increases with a decrease in the Kemeny constant of the surveillance policy; the policy with the lowest Kemeny constant catches the most intruders. Inspecting the $P(x^{\text{non-rev}})$ policy in Fig. 6, we see that a strong level of (clockwise) directionality is achieved (note that the Hamiltonian cycle is the extreme case of directionality), which the $P(x^{\text{rev}})$ policy cannot obtain. This directionality feature is crucial for our approach to obtain lower MFPTs and, thus, higher intruder detection rates.
@@ VII-2 Random support
[…]
@@ VIII Conclusion and future work
To illustrate why standard SPSA does not work, let us consider the following Markov chain on the complete graph $\mathscr{G}$ with $N=3$ nodes and stochastic matrix
@@
[…]
@@
Observe that if the weighted graph $(\mathscr{G},x)$, where $x\in\mathcal{X}^{\mathrm{sym}}$, is strongly connected, the Markov chain $P(x)$ via (3) satisfies (R), i.e., $P(x)\in\mathcal{P}^{\mathrm{rev}}$ [10]. Moreover, it has stationary distribution $\pi$, where for all $i\in V$
@@
This uniqueness implies that $x$ is a unique representation for $P$ via $P(x)$. Moreover, since $x$ is strongly connected as follows from the irreducibility of $P\in\mathcal{P}^{\mathrm{rev}}$, the mapping $P(x)$ is a bijection from $\mathcal{X}^{(\mathrm{sym},1)}\cap\{x\in\mathcal{X}^{\mathrm{sym}}:(\mathscr{G},x)\text{ is strongly connected}\}$ to $\mathcal{P}^{\mathrm{rev}}$.
@@
The optimization problem (17) has been studied in detailed in [24], in which the authors proved that the optimal solution $\hat{P}$ is a Hamiltonian cycle, see Theorem 2 therein. The same paper also shows that $\sum_{i\neq j}M_{ij}=\frac{N(N-1)}{2}$ for all $i$, from which it easily follows that
@@
Consider the family of complete undirected graphs $(\mathscr{G},x)$ with $x\in\mathcal{X}^{\mathrm{sym}}$. The complete graph with $N$ nodes, with $x_{\ell}=1$, for all $\ell\in E$, has been shown to have total effective graph resistance $R_{tot}(x)=N-1$, see [7]. Using this fact in combination with (8) and (9) provides the lower bound in (20), that is
@@
The upper bound in (20) follows from the fact that the Hamiltonian graph with maximal $R_{tot}(x)$ is a circular graph (as adding edges to $\mathscr{G}$ cannot increase $R_{tot}(x)$, [10]). In [44], it has been shown that for a circular graph $(\mathscr{G},x)$ of size $N$, $R_{tot}(x)=\frac{N^{3}-N}{12}$, where $x_{\ell}=1$ if $\ell\in E$ is part of the circle and $x_{\ell}=0$ otherwise. Since $(\mathscr{G},x)$ is a circular graph, we have $\sum_{\ell\in E}x_{\ell}=N$. Rewriting (8) yields: